Felipe purchased a prepaid phone card for $35.50. Calls cost 7 cents a minute using this card. The credit, C(in dollars), left on the card after it is used for x minutes of calls is given by the following function. C(x)-35.50-0.07x . How much credit is left on the card after Felipe uses it for 40 minutes of calls?
Answer by Cromlix(4381)

Hi there,
I think there maybe a typo
on your formula.
Think it should read:-
C(x) = 35.50 - 0.07x
After Felipe has used it for
40 minutes:-
C(x) = 35.50 - (0.07 x 40)
C(x) = 35.50 - 2.8
C(x) = $32.7 left.
Hope this helps :-)